2a cubed plus (2a) cubed
1
Expert's answer
2013-03-25T10:34:58-0400
We have the following expression:

2a³ + (2a)³.

Let's transform it. We'll open the brackets at first:

2a³ + (2a)³ = 2a³ + 2³a³ = 2a³ + 8a³.

We've got two similar terms, so let's add them:

2a³ + 8a³ = (2+8)a³ = 10a³.

So, 2a³ + (2a)³ = 10a³.